Farmer Tariq's rectangular field is 40 meters long. If the area of the field is 1,000 square meters, how wide, in meters, is the field?
Answer by haileytucki(390)

L*W=area
40*x=1000
Multiply 40 by x to get 40x.
40x=1000
Divide each term in the equation by 40.
(40x)/(40)=(1000)/(40)
Simplify the left-hand side of the equation by canceling the common factors.
x=(1000)/(40)
Simplify the right-hand side of the equation by simplifying each term.
x=25 meters